### v1

@@ Abstract
Multicategory remote object counting is a fundamental task in computer vision, aimed at accurately estimating the number of objects of various categories in remote images. Existing methods rely on CNNs and Transformers, but CNNs struggle to capture global dependencies, and Transformers are computationally expensive, which limits their effectiveness in remote applications. Recently, Mamba has emerged as a promising solution in the field of computer vision, offering a linear complexity for modeling global dependencies. To this end, we propose Mamba-MOC, a mamba-based network designed for multi-category remote object counting, which represents the first application of Mamba to remote sensing object counting. Specifically, we propose a cross-scale interaction module to facilitate the deep integration of hierarchical features. Then we design a context state space model to capture both global and local contextual information and provide local neighborhood information during the scan process. Experimental results in large-scale realistic scenarios demonstrate that our proposed method achieves state-of-the-art performance compared with some mainstream counting algorithms. The code will be available at https://github.com/lp-094/Mamba-MOC.
@@ I Introduction
Conventional methods for remote sensing object counting predominantly rely on Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs). CNN-based approaches are widely utilized due to their translational invariance and linear computational complexity. However, the fixed-size local connections inherent in CNNs constrain their receptive fields, limiting the ability to effectively capture long-range dependencies and hindering the modeling of global context. In contrast, ViT-based approaches exploit long-range dependencies, which are important for visual tasks [4, 5] and dynamic weight allocation to enhance the modeling of visual information. However, the self-attention mechanism inherent in ViTs introduces quadratic complexity with respect to input size, which results in a significant increase in computational cost, especially when handling large-scale data.
@@ II Preliminaries
$\begin{split}\mathbf{\overline{A}}=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{AB}},...,\mathbf{C}{\mathbf{\overline{A}}}^{L-1}\mathbf{\overline{B}}),y=x\ast\overline{\mathbf{K}}\end{split}$ (4)
@@ III-C Context State Space Model
Inspired by [15], the output matrix $\mathbf{C}$ in Equation 1 functions as a query in the attention mechanism. From this observation, we derive two key insights: (1) Through local 2D modeling, $\mathbf{C}$ effectively “queries” pixels within the local context, addressing the inherent limitations of causal scanning in the Mamba model when applied to 2D images. (2) By emphasizing local context, the model focuses on fine-grained details while simultaneously integrating global context through global scanning, which improves the extraction of target features.
Based on this concept, we propose a Context State Space Model (CSSM) in a straightforward manner. As shown in Fig. 3, the proposed CSSM incorporates multi-scale local information, allowing the model to learn and capture correlations between local neighbourhoods within the image. This method implicitly mitigates the inherent limitations of the SSM, which compromise the preservation of spatial structure. At the same time, it integrates both local and global contextual information, improving the perception of counting targets. Specifically, given an input feature map $F_{c}\in\mathbb{R}^{H\times W\times C}$, we first capture multi-scale contextual features $F_{ms}\in\mathbb{R}^{H\times W\times D}$ by applying multiple convolutional operations, from which $P\in\mathbb{R}^{HW\times D}$ is obtained through a flattening step. Simultaneously, a $1\times 1$ convolution is applied for linear projection to yield $D\in\mathbb{R}^{H\times W\times C}$, as expressed in the following equation:
$\begin{split}F_{ms}&=CR_{3\times 3,d=1}(F_{c})+CR_{3\times 3,d=2}(F_{c}),\\ P&=Flatten(F_{ms}),\\ D&=Conv_{1\times 1}(F_{ms})\end{split}$ (7)
@@
Then, we incorporate $P$ into $C$ throgh residual addition and incorporate $D$ into $y(t)$ to formulate the context state space model:
$\begin{split}h^{{}^{\prime}}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\\ y(t)=(\mathbf{C}+\mathbf{P})h(t)+\mathbf{D}\end{split}$ (8)

### v2

@@ Abstract
Multicategory remote object counting is a fundamental task in computer vision, aimed at accurately estimating the number of objects of various categories in remote sensing images. Existing methods rely on CNNs and Transformers, but CNNs struggle to capture global dependencies, and Transformers are computationally expensive, which limits their effectiveness in remote applications. Recently, Mamba has emerged as a promising solution in the field of computer vision, offering a linear complexity for modeling global dependencies. To this end, we propose Mamba-MOC, a mamba-based network designed for multi-category remote object counting, which represents the first application of Mamba to remote sensing object counting. Specifically, we propose a cross-scale interaction module to facilitate the deep integration of hierarchical features. Then we design a context state space model to capture both global and local contextual information and provide local neighborhood information during the scan process. Experimental results in large-scale realistic scenarios demonstrate that our proposed method achieves state-of-the-art performance compared with some mainstream counting algorithms. The code will be available at https://github.com/lp-094/Mamba-MOC.
@@ I Introduction
Conventional methods for remote sensing object counting predominantly rely on Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs). CNN-based approaches are widely utilized due to their translational invariance and linear computational complexity. However, the fixed-size local connections inherent in CNNs constrain their receptive fields, limiting the ability to effectively capture long-range dependencies and hindering the modeling of global context. In contrast, ViT-based approaches exploit long-range dependencies, which are important for visual tasks [4] and dynamic weight allocation to enhance the modeling of visual information. However, the self-attention mechanism inherent in ViTs introduces quadratic complexity with respect to input size, which results in a significant increase in computational cost, especially when handling large-scale data.
@@ II Preliminaries
$\begin{split}\mathbf{\overline{K}}=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{AB}},...,\mathbf{C}{\mathbf{\overline{A}}}^{L-1}\mathbf{\overline{B}}),y=x\ast\overline{\mathbf{K}}\end{split}$ (4)
@@ III-C Context State Space Model
Inspired by [14], the output matrix $\mathbf{C}$ in Equation 1 functions as a query in the attention mechanism. This analysis yields two critical findings: (1) Through local 2D modeling, $\mathbf{C}$ effectively “queries” pixels within the local context, addressing the inherent limitations of causal scanning in the Mamba model when applied to 2D images. (2) By emphasizing local context, the model focuses on fine-grained details while simultaneously integrating global context through global scanning, which improves the extraction of target features.
Based on this concept, we propose a Context State Space Model (CSSM) in a straightforward manner. As shown in Fig. 3, the proposed CSSM incorporates multi-scale local information, allowing the model to learn and capture correlations between local neighbourhoods within the image. This method implicitly mitigates the inherent limitations of the SSM, which compromise the preservation of spatial structure. At the same time, it integrates both local and global contextual information, improving the perception of counting targets. Specifically, given an input feature map $F_{c}\in\mathbb{R}^{H\times W\times C}$, we first capture multi-scale contextual features $F_{ms}\in\mathbb{R}^{H\times W\times D}$ by applying multiple convolutional operations, from which local query $Q_{l}\in\mathbb{R}^{HW\times D}$ is obtained through a flattening step. Simultaneously, a $1\times 1$ convolution is applied for linear projection to yield local feature $F_{l}\in\mathbb{R}^{H\times W\times C}$, as expressed in the following equation:
$\begin{split}F_{ms}&=CR_{3\times 3,d=1}(F_{c})+CR_{3\times 3,d=2}(F_{c}),\\ Q_{l}&=Flatten(F_{ms}),\\ F_{l}&=Conv_{1\times 1}(F_{ms})\end{split}$ (7)
@@
Then, we incorporate $Q_{l}$ into $C$ throgh residual addition and incorporate $F_{l}$ into $y(t)$ to formulate the context state space model:
$\begin{split}h^{{}^{\prime}}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\\ y(t)=(\mathbf{C}+\mathbf{Q_{l}})h(t)+\mathbf{F_{l}}\end{split}$ (8)